# Fixation and extinction in time-fluctuating spatially structured metapopulations

### v2

@@ III.2 Extinction-dominated dynamics
In the extinction-dominated regime $\psi<1$ with $m\ll 1$ (slow migration), we do not expect any deme invasions in a time $\tau_{E}$, enabling us to adopt a suitable coarse-grained description. Deme invasions being negligible under slow migration, the timescale of extinction dynamics is much shorter than that of $M/W$ competition, and site extinction dominates over deme invasion; see Fig. 1(e). The dynamics in this extinction-dominated regime is hence governed by the random extinction of demes, regardless of their type. Following Ref. [59], in the realm of the coarse-grained description outlined in Appendix C, demes are regarded as being either occupied (by either $W$ or $M$ cells) or empty. Deme extinction occurs randomly while empty demes may be recolonised by migrations from occupied neighbouring demes, which occur rarely in this regime. Since spatial structure may only influence the dynamics via migrations, the coarse-grained dynamics is largely independent of the spatial structure in this regime where $m\ll 1$, and can be represented by a birth-death process for the number of occupied demes, assuming that the site extinction and recolonisation occur instantaneously [59]; see Appendix C. In this coarse-grained representation, when all demes are initially occupied ($n=K$), the metapopulation mean extinction time (mMET), denoted by $\theta_{E}$, is given by Eq. (39) in Appendix C. When $\Omega\gg 1$ and $\psi<1$, the leading contribution to the mMET arises from the term $i=n$ in the innermost sum, yielding
@@
where $\gamma_{\text{EM}}\approx 0.577$... is the Euler-Mascheroni constant. Eq. (15) gives a good approximation of the mMET as reported in Fig. 8. As $\psi$ increases with $K$, at the upper-limit of the extinction-dominated limit (where $\psi$ approaches $1$), the mMET grows almost exponentially with $K$ and logarithmically with $\Omega$, $\theta_{E}\approx e^{K}\ln{(\Omega)}/K$. Eq. (39) thus predicts a rapid growth of the mMET when $\psi\gtrsim 1$, as shown in Fig. 2(f) where simulation results are found to be in good agreement with Eq. (39).
III.3 Intermediate dynamics
In the intermediate regime $\psi\gtrsim 1$ with $m\ll 1$ (slow migration), there is a crossover between the extinction and competition-dominated dynamics that we can characterise at a coarse-grained level. At fixed migration rate $m$, this crossover regime occurs in the range $\ln{(1/m)}\lesssim K\lesssim\ln{(\Omega K/m)}$; see Appendix F. Furthermore, the metapopulation is partially occupied for $\ln{(1/m)}\lesssim K\lesssim\ln{(\Omega/m)}$ corresponding to $1<\psi<\Omega-1$. These bounds are illustrated by the vertical lines in Fig. 11.
When $\psi\gtrsim 1$, migration and deme extinction occur on the timescale $\tau_{E}$. In the long run, deme recolonisations and extinctions balance each other, yielding a dynamical equilibrium consisting of $\Omega_{\rm occ}=\Omega(1-1/\psi)$ occupied demes and $\Omega-\Omega_{\rm occ}$ empty demes; see Fig. 9. In this regime, the three-state coarse-grained description of the dynamical equilibrium consists of a random mixture of empty demes, and occupied $W$ and $M$ demes; see Fig. 1(d). After a mean time $\theta_{\rm int}^{\rm G}$ the metapopulation reaches the dynamical equilibrium consisting of a fraction $1-\Omega_{\rm occ}/\Omega=1/\psi$ empty demes, and the remaining demes are all of either type $M$ or $W$ with probability $\phi^{G}_{\rm int}$ and $1-\phi_{\rm int}^{\rm G}$, respectively. The dynamical equilibrium is thus defined by the quantities $\psi$, given by Eq. (8), and $\phi_{\rm int}^{\rm G}$ which is the probability that mutants $M$ take over the $\Omega_{\rm occ}=\Omega(1-1/\psi)$ occupied demes, where the unconditional mean fixation time is given by $\theta_{\rm int}^{\rm G}$. These quantities are derived and discussed in Appendix F for a metapopulation on an arbitrary regular lattice consisting initially of a single $M$ deme; see Eq. (44) and Fig. 11.
@@ IV Time-fluctuating environments
Microbial populations generally evolve in time-varying environments, and are often subject to conditions changing suddenly and drastically, e.g. experiencing cycles of harsh and mild environmental states [73, 20, 74, 30, 75, 76, 77, 78, 79, 28, 80, 81, 82, 83, 84], see Fig. 3. These variations cause fluctuations often associated with population bottlenecks, arising when the deme size is drastically reduced, e.g. due to nutrient scarcity or exposure to toxins [78, 28, 29, 27, 31]. Here, environmental variability is encoded in the time-fluctuating carrying capacity $K(t)$ of Eq. (6) driven by the DMN $\xi(t)\in\{-1,1\}$ [68, 69, 70, 16, 17, 18, 23, 20, 19, 21, 22, 24, 32]; see Sec. II. Since the dynamics of the deme size $n$ occurs on a timescale of order $1$ (see Appendix A.2), the variable $n$ tracks $K(t)$ and experiences a bottleneck whenever the carrying capacity switches from $K_{+}$ to $K_{-}$ at a rate $\nu\lesssim 1$, see Fig. 3(a,right) [16, 17, 18].
@@ IV.1 Weak bottlenecks: $\psi(m,K_{-})\gg 1$
Figure 4: Fixation probability $\Phi^{\rm G}$ and mean fixation time $\Theta^{\rm G}$ against switching rate $\nu$ for various parameters. Each panel shows $\Phi^{\rm G}$ vs. $\nu$ (left) and $\Theta^{\rm G}$ vs. $\nu$ (right). Markers show simulation results and lines are predictions of Eq. (26). (a,b) $\Phi^{\rm clique}(\nu)$ and $\Theta^{\rm clique}(\nu)$ for a clique metapopulation and different values of $m$ in (a) and $s$ in (b). (a) $m=10^{-5}$ (red), $m=10^{-4}$ (blue), $m=10^{-3}$ (yellow), and $s=0.01$. (b) $s=10^{-3}$ (red), $s=10^{-2}$ (blue), $s=10^{-1}$ (yellow), and $m=10^{-4}$. Dashed black lines are guides to the eye showing $\Phi_{0,\infty}$ in (a,left) and $\Theta_{0,\infty}$ in (a,right); see text. Other parameters are $\Omega=16$, $K_{+}=200$, and $K_{-}=20$. (c) $\Phi^{\rm G}(\nu)$ and $\Theta^{\rm G}(\nu)$ for clique (red, crosses), cycle (blue, circles), and grid (yellow, triangles) metapopulations. Other parameters are $\Omega=16$, $K_{+}=200$, $K_{-}=20$, $s=0.01$, $m=10^{-4}$. (d) $\Phi^{\rm clique}(\nu)$ and $\Theta^{\rm clique}(\nu)$ for a clique metapopulation with $K_{+}=200$ (red), $K_{+}=500$ (blue), and $K_{+}=1000$ (yellow). Deviations occur for $\Theta$ with $K_{+}=1000$ since the slow-migration condition is not satisfied in the mild environment. Other parameters are $\Omega=16$, $K_{-}=20$, and $s=0.01$, $m=10^{-4}$. In all examples, there is initially a single $M$ deme and $\Omega-1$ others of type $W$; see text.
When $K_{+}>K_{-}$ with $\psi(m,K_{-})\gg 1$, each deme experiences a weak bottleneck at an average frequency $\nu/2$ when $\nu\lesssim 1$ [16, 17, 18]; see Fig. 3(a,top). The condition $\psi(m,K_{-})\approx me^{K_{-}}\gg 1$ ensures that deme extinction can be neglected, with metapopulation dynamics dominated by $M/W$ competition. The metapopulation fate can thus be captured by a two-state coarse-grained description similar to that of Sec. III.1. Since the deme size in the environment $\xi$, and hence the number of migrating cells, varies with $K(t)$, it is useful to introduce the long-term average deme size in environmental state $\xi=\pm 1$ under switching rate $\nu$ denoted by $\mathcal{N}_{\xi}(\nu)\equiv\mathcal{N}_{\pm}(\nu)$; see below.
@@
Under intermediate switching rate, when $\nu\lesssim 1$, the coupling of demographic and environmental fluctuations plays a key role, while cell migration depends on the deme size that in turn varies with the environmental state. Here, the metapopulation dynamics cannot be directly related to its static counterpart. The average deme size $\mathcal{N}_{\xi}(\nu)$ depends non-trivially on $\nu$ and $\xi$, and generally needs to be computed by sampling long-time simulations. However, analytical progress can be made by approximating the distribution of the size $n$ of an isolated deme in the environmental state $\xi$ by the joint probability density $p_{\xi}(n;\nu)$ of the piecewise deterministic Markov process (PDMP), where $n$ and $\xi$ are variables, and $\nu$ a parameter, obtained by ignoring demographic fluctuations [85, 86, 16, 17]; see Appendix A.2:
@@
With Eq. (21), by dropping all explicit dependence except on $i$ and $\xi$, we obtain the $M$ fixation probability starting from the environmental state $\xi$ with $i$ mutant demes on a graph $\rm G$, denoted by $\Phi_{i,\xi}^{{\rm G}}(\nu,m)$, as the solution of the following $\nu$-dependent first-step analysis equation:
@@
subject to the boundary conditions $\Phi_{0,\xi}^{{\rm G}}=0$ and $\Phi_{\Omega,\xi}^{{\rm G}}=1$. The metapopulation uMFT starting from the same initial conditions, denoted by $\Theta_{i,\xi}^{{\rm G}}$, satisfies a similar equation:
@@
$\Phi^{\rm G}(\nu,m)=\Phi_{1}^{\rm G}(\nu,m)\text{ and }\Theta^{\rm G}(\nu,m)=\Theta_{1}^{\rm G}(\nu,m),$ (26)
@@
Eq. (26) are the expressions of the $M$ fixation probability and metapopulation uMFT in the realm of the combined coarse-grained and PDMP description. This approach is valid under slow migration rate ($m\ll 1$) and weak selection strength ($s\ll 1$) for the assumption $n\approx\mathcal{N}_{\xi}(\nu)$ to hold at each invasion; see Appendix A.2. In Fig. 4, the comparison of the predictions of Eq. (26) with the simulation results of the full model on the regular graphs ${\rm G}=\{{\rm clique,cycle,grid}\}$ shows that Eq. (26) captures well the dependence of $\Phi^{\rm G}$ and $\Theta^{\rm G}$ on $\nu$, $m$, $s$ and $K_{+}$. In particular, Eq. (26) reproduces on all $\rm G$} the non-monotonic $\nu$-dependence of $\Phi^{\rm G}$ and $\Theta^{\rm G}$ (when it exhibits this feature), as well as their behaviour when $\nu\to 0,\infty$ given by Eqs. (16) and (18).
@@
A striking feature of $\Phi^{\rm G}$ and $\Theta^{\rm G}$ is their dependence on spatial migration. In Fig. 4(a,left), we indeed find that simulation data for $\Phi^{\rm clique}(\nu,m)$ vary noticeably with $m$ in the range $\nu\in[10^{-4},10^{-1}]$. These deviations, of up to $20\%$, exceed the error bars and are reasonably well captured by Eq. (26). In Fig. 4(c,left), we notice that both simulation results and predictions of Eq. (26) for $\Phi^{\rm G}(\nu,m)$ differ slightly for each graph $\rm G$}, whereas Fig. 4(c,right) shows that $\Theta^{\rm G}$ clearly depends on the spatial structure. The explicit dependence of the fixation probability on migration and spatial structure is in stark contrast with the result Eq. (14) obtained in static environments, and is therefore a signature of the eco-evolutionary dynamics in time-fluctuating environments. As shown in Appendix G, the correspondence demonstrated in Ref. [41] between $\Phi^{\rm G}$ and the fixation probability of a random walk for the number $i=0,1,\dots,\Omega$ of mutant demes with hopping probabilities independent of $m$ and $\rm G$}, and absorbing states $0,\Omega$, breaks down in time-varying environments. This leads to the dependence of $\Phi^{\rm G}$ and $\Theta^{\rm G}$ on $m$ and $\rm G$} in time-switching environments.
@@
Another distinctive feature of $\Phi^{\rm G}$ and $\Theta^{\rm G}$ is their non-monotonic $\nu$-dependence when the other parameters ($s$, $m$, $K_{\pm}$, $\Omega$) are kept fixed. In particular, Fig. 4 shows that $\Phi^{\rm G}$ may exhibit a sharp peak in the regime of intermediate $\nu\in[10^{-4},10^{-1}]$ that is well captured by Eq. (26). These results are in marked contrast with their counterparts in a single deme, which vary monotonically with $\nu$ [16, 17, 19]. The non-monotonic $\nu$-dependence of $\Phi^{\rm G}$ and $\Theta^{\rm G}$ is therefore an inherent effect of spatial migration. Intuitively, this behaviour stems, on the one hand, from more $M$ invasions occurring when the deme size is as close as possible to $n\approx K_{+}$; see Eq. (21). On the other hand, the average deme size is a decreasing function of $\nu$ [16, 17, 18]; see Appendix A.2. Therefore, optimising the probability of $M$ fixation requires two considerations: the environment should avoid remaining stuck in the harsh environment for too long, which can happen with a probability close to $1/2$ when $\nu\ll 1$; and the environment should not switch too frequently (i.e. $\nu\gg 1$), as this reduces the effective deme size ($n\approx{\cal K}$), which can be significantly smaller than $K_{+}$. Hence, the best conditions for the fixation of $M$ are for a range of $\nu$ in the intermediate regime. Since the uMFT is longer in the harsh than in the mild environment (where there are fewer migration events; see Fig. 2(b)), a similar reasoning leads to a minimum mean fixation time for $\nu$ in the intermediate regime; see Fig. 4(b,d).

### v3

@@ III.2 Extinction-dominated dynamics
In the extinction-dominated regime $\psi<1$ with $mK<1$ (slow migration), we do not expect any deme invasions in a time $\tau_{E}$, enabling us to adopt a suitable coarse-grained description. Deme invasions being negligible under slow migration, the timescale of extinction dynamics is much shorter than that of $M/W$ competition, and site extinction dominates over deme invasion; see Fig. 1(e). The dynamics in this extinction-dominated regime is hence governed by the random extinction of demes, regardless of their type. Following Ref. [59], in the realm of the coarse-grained description outlined in Appendix C, demes are regarded as being either occupied (by either $W$ or $M$ cells) or empty. Deme extinction occurs randomly while empty demes may be recolonised by migrations from occupied neighbouring demes, which occur rarely in this regime. Since spatial structure may only influence the dynamics via migrations, the coarse-grained dynamics is largely independent of the spatial structure in this regime where $mK<1$, and can be represented by a birth-death process for the number of occupied demes, assuming that the site extinction and recolonisation occur instantaneously [59]; see Appendix C. In this coarse-grained representation, when all demes are initially occupied ($n=K$), the metapopulation mean extinction time (mMET), denoted by $\theta_{E}$, is given by Eq. (39) in Appendix C. When $\Omega\gg 1$ and $\psi<1$, the leading contribution to the mMET arises from the term $i=n$ in the innermost sum, yielding
@@
where $\gamma_{\text{EM}}\approx 0.577$... is the Euler-Mascheroni constant. Eq. (15) gives a good approximation of the mMET as reported in Fig. 8. As $\psi$ increases with $K$, at the upper-limit of the extinction-dominated limit (where $\psi$ approaches $1$), the mMET grows almost exponentially with $K$ and logarithmically with $\Omega$, $\theta_{E}\approx e^{K}\ln{(\Omega)}/K$. Eq. (39) thus predicts a rapid growth of the mMET when $\psi\gtrsim 1$, as shown in Fig. 2(b,bottom) where simulation results are found to be in good agreement with Eq. (39).
When $\psi\gtrsim 1$, the competition- and extinction-dominated dynamics are separated by an intermediate crossover regime analysed in detail in Appendix F.
@@ IV Time-fluctuating environments
Microbial populations generally evolve in time-varying environments, and are often subject to conditions changing suddenly and drastically, e.g. experiencing cycles of harsh and mild environmental states [73, 20, 74, 30, 75, 76, 77, 78, 79, 28, 80, 81, 82, 83, 84], see Fig. 3. These variations cause fluctuations often associated with population bottlenecks, arising when the deme size is drastically reduced, e.g. due to nutrient scarcity or exposure to toxins [78, 28, 29, 27, 31]. Here, environmental variability is encoded in the time-fluctuating carrying capacity $K(t)$ of Eq. (6) driven by the DMN $\xi(t)\in\{-1,1\}$ [68, 69, 70, 16, 17, 18, 23, 20, 19, 21, 22, 24, 32]; see Sec. II. Since the dynamics of the deme size $n$ occurs on a timescale of order $1$ (see Appendix A.2), the variable $n$ tracks $K(t)$ and experiences a bottleneck whenever the carrying capacity switches from $K_{+}$ to $K_{-}$ at a rate $\nu\lesssim 1$; see Fig. 3(a,right) [16, 17, 18].
@@ IV.1 Weak bottlenecks: $\psi(m,K_{-})\gg 1$
Figure 4: Fixation probability $\Phi^{\rm G}$ and mean fixation time $\Theta^{\rm G}$ against switching rate $\nu$ for various parameters. Each panel shows $\Phi^{\rm G}$ vs. $\nu$ (top) and $\Theta^{\rm G}$ vs. $\nu$ (bottom). Markers show simulation results and lines are predictions of Eq. (26). (a,b) $\Phi^{\rm clique}(\nu)$ and $\Theta^{\rm clique}(\nu)$ for a clique metapopulation and different values of $m$ in (a) and $s$ in (b). (a) $m=10^{-5}$ (red), $m=10^{-4}$ (blue), $m=10^{-3}$ (yellow), and $s=0.01$. (b) $s=10^{-3}$ (red), $s=10^{-2}$ (blue), $s=10^{-1}$ (yellow), and $m=10^{-4}$. Dashed black lines are guides to the eye showing $\Phi_{0,\infty}$ in (a,top) and $\Theta_{0,\infty}$ in (a,bottom); see text. Other parameters are $\Omega=16$, $K_{+}=200$, and $K_{-}=20$. (c) $\Phi^{\rm G}(\nu)$ and $\Theta^{\rm G}(\nu)$ for clique (red, crosses), cycle (blue, circles), and grid (yellow, triangles) metapopulations. Other parameters are $\Omega=16$, $K_{+}=200$, $K_{-}=20$, $s=0.01$, $m=10^{-4}$. (d) $\Phi^{\rm clique}(\nu)$ and $\Theta^{\rm clique}(\nu)$ for a clique metapopulation with $K_{+}=200$ (red), $K_{+}=500$ (blue), and $K_{+}=1000$ (yellow). Deviations occur for $\Theta$ with $K_{+}=1000$ since the slow-migration condition is not satisfied in the mild environment. Other parameters are $\Omega=16$, $K_{-}=20$, and $s=0.01$, $m=10^{-4}$. In all examples, there is initially a single $M$ deme and $\Omega-1$ others of type $W$; see text.
When $K_{+}>K_{-}$ with $\psi(m,K_{-})\gg 1$, each deme experiences a weak bottleneck at an average frequency $\nu/2$ when $\nu\lesssim 1$ [16, 17, 18]; see Fig. 3(a,top). The condition $\psi(m,K_{-})\approx me^{K_{-}}\gg 1$ ensures that deme extinction can be neglected, with metapopulation dynamics dominated by $M/W$ competition. The metapopulation fate can thus be captured by a two-state coarse-grained description similar to that of Sec. III.1. Since the deme size in the environment $\xi$, and thus the number of migrating cells, varies with $K(t)$, it is useful to introduce the long-time average deme size in environmental state $\xi=\pm 1$ under switching rate $\nu$ denoted by $\mathcal{N}_{\xi}(\nu)\equiv\mathcal{N}_{\pm}(\nu)$; see below.
@@
Under intermediate switching rate, when $\nu\lesssim 1$, the coupling of demographic and environmental fluctuations plays a key role, while cell migration depends on the deme size that in turn varies with the environmental state. Here, the metapopulation dynamics cannot be directly related to its static counterpart. The average deme size $\mathcal{N}_{\xi}(\nu)$ depends non-trivially on $\nu$ and $\xi$, and generally needs to be computed by sampling long-time simulations. However, analytical progress can be made by approximating the distribution of the size $n$ of an isolated deme in the environmental state $\xi$ by the joint probability density $p_{\xi}(n;\nu)$ of the piecewise deterministic Markov process (PDMP), where $n$ and $\xi$ are variables, and $\nu$ a parameter, obtained by ignoring demographic fluctuations [85, 86, 16, 17] (see Appendix A.2):
@@
With Eq. (21), by dropping all explicit dependence except on $i$ and $\xi$, we obtain the $M$ fixation probability starting from the environmental state $\xi$ with $i$ mutant demes on a graph $\rm G$, denoted by $\Phi_{i,\xi}^{{\rm G}}(\nu,m)$, as the solution of the $\nu$-dependent first-step analysis equation
@@
subject to the boundary conditions $\Phi_{0,\xi}^{{\rm G}}=0$ and $\Phi_{\Omega,\xi}^{{\rm G}}=1$. The metapopulation uMFT starting from the same initial conditions, denoted by $\Theta_{i,\xi}^{{\rm G}}$, similarly satisfies
@@
$\Phi^{\rm G}(\nu,m)=\Phi_{1}^{\rm G}(\nu,m)\text{ and }\Theta^{\rm G}(\nu,m)=\Theta_{1}^{\rm G}(\nu,m).$ (26)
@@
Eq. (26) are the expressions of the $M$ fixation probability and metapopulation uMFT in the realm of the combined coarse-grained and PDMP description. This approach is valid under slow migration rate ($mK<1$) and weak selection strength ($s\ll 1$) for the assumption $n\approx\mathcal{N}_{\xi}(\nu)$ to hold at each invasion; see Appendix A.2. In Fig. 4, the comparison of the predictions of Eq. (26) with the simulation results of the full model on the regular graphs ${\rm G}=\{{\rm clique,cycle,grid}\}$ shows that Eq. (26) captures well the dependence of $\Phi^{\rm G}$ and $\Theta^{\rm G}$ on $\nu$, $m$, $s$ and $K_{+}$. In particular, Eq. (26) reproduces on all $\rm G$} the non-monotonic $\nu$-dependence of $\Phi^{\rm G}$ and $\Theta^{\rm G}$ (when it exhibits this feature), as well as their behaviour when $\nu\to 0,\infty$ given by Eqs. (16) and (18).
@@
A striking feature of $\Phi^{\rm G}$ and $\Theta^{\rm G}$ is their dependence on spatial migration. In Fig. 4(a,top), we indeed find that simulation data for $\Phi^{\rm clique}(\nu,m)$ vary noticeably with $m$ in the range $\nu\in[10^{-4},10^{-1}]$. These deviations, of up to $20\%$, exceed the error bars and are reasonably well captured by Eq. (26). In Fig. 4(c,top), we notice that both simulation results and predictions of Eq. (26) for $\Phi^{\rm G}(\nu,m)$ differ slightly for each graph $\rm G$}, whereas Fig. 4(c,bottom) shows that $\Theta^{\rm G}$ clearly depends on the spatial structure. The explicit dependence of the fixation probability on migration and spatial structure is in stark contrast with the result Eq. (14) obtained in static environments, and is therefore a signature of the eco-evolutionary dynamics in time-fluctuating environments. As shown in Appendix G, the correspondence demonstrated in Ref. [41] between $\Phi^{\rm G}$ and the fixation probability of a random walk for the number $i=0,1,\dots,\Omega$ of mutant demes with hopping probabilities independent of $m$ and $\rm G$}, and absorbing states $0,\Omega$, breaks down in time-varying environments. This leads to the dependence of $\Phi^{\rm G}$ and $\Theta^{\rm G}$ on $m$ and $\rm G$} in time-switching environments.
@@
Another distinctive feature of $\Phi^{\rm G}$ and $\Theta^{\rm G}$ is their non-monotonic $\nu$-dependence when the other parameters ($s$, $m$, $K_{\pm}$, $\Omega$) are kept fixed. In particular, Fig. 4 shows that $\Phi^{\rm G}$ may exhibit a sharp peak in the regime of intermediate $\nu\in[10^{-4},10^{-1}]$ that is well captured by Eq. (26). These results are in marked contrast with their counterparts in a single deme, which vary monotonically with $\nu$ [16, 17, 19]. The non-monotonic $\nu$-dependence of $\Phi^{\rm G}$ and $\Theta^{\rm G}$ is therefore an inherent effect of spatial migration. Intuitively, this behaviour stems, on the one hand, from more $M$ invasions occurring when the deme size is as close as possible to $n\approx K_{+}$; see Eq. (21). On the other hand, the average deme size is a decreasing function of $\nu$ [16, 17, 18]; see Appendix A.2. Therefore, optimising the probability of $M$ fixation requires two considerations: the environment should avoid remaining stuck in the harsh environment for too long, which can happen with a probability close to $1/2$ when $\nu\ll 1$; and the environment should not switch too frequently (i.e. $\nu\gg 1$), as this reduces the effective deme size ($n\approx{\cal K}$), which can be significantly smaller than $K_{+}$. Hence, the best conditions for the fixation of $M$ are for a range of $\nu$ in the intermediate regime. Since the uMFT is longer in the harsh than in the mild environment (where there are fewer migration events; see Fig. 2(a,bottom left)), a similar reasoning leads to a minimum mean fixation time for $\nu$ in the intermediate regime; see Fig. 4(b,d).